[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Oct 05, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Oct 6 03:14:27 UTC 2017


HSU Shift Log for Thu Oct 05 2017 23:14:20 GMT-0400 (EDT)

Shift: R

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: Some issues with creating the watch/warning graphic on  
ridges and hurrback. The vgf would not open on those computers. Dave Z.  
suggested that they might need to be rebooted.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171006/35f04a71/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list